# GS-LiDAR: Generating Realistic LiDAR Point Clouds with Panoramic Gaussian Splatting

### v1

@@ Abstract
Figure 1: GS-LiDAR achieves superior LiDAR simulation quality for novel view synthesis while maintaining fast training and rendering speed. For more results on static scene, dynamic scene, lane shifting and novel trajectory, please refer to the video in our supplementary material.
@@ 3.4 LiDAR rendering
Figure 5: Comparison of 3D LiDAR point cloud. GS-LiDAR produces a more cohesive LiDAR point cloud compared to LiDAR-NeRF (Tao et al., 2023) and LiDAR4D (Zheng et al., 2024). For more results, please refer to the video in our supplementary material.
@@ 4.1 Experiment setup
We conduct extensive experiments on both dynamic and static scenes using the KITTI-360 (Liao et al., 2022) and nuScenes (Caesar et al., 2020) datasets, with the dynamic scenes featuring a significant number of moving vehicles. The KITTI-360 dataset employs a 64-beam LiDAR with a vertical field of view (FOV) of 26.4 degrees and an acquisition frequency of 10Hz. Following LiDAR4D (Zheng et al., 2024), we select 51 consecutive frames as a single scene and hold out 4 samples at 10-frame intervals for novel view synthesis (NVS) evaluation. For the nuScenes dataset, the LiDAR system uses 32 beams with a 40-degree vertical FOV and a 20Hz acquisition frequency. To ensure consistency with KITTI-360, we maintain a sampling frequency of 10Hz.
Figure 6: Comparison of the rendered depth map with competitors. For more results, please refer to the video in our supplementary material.
@@ 4.2 Evaluation on static scenes
Figure 8: Comparison of w/ and w/o ray-splat intersection.
@@ 4.4 Ablation study
We provide quantitative ablation studies on various components of GS-LiDAR in Table 4. As shown in Figure 8, we find that the use of ray-splat intersection enhances the capability of GS-LiDAR in modeling the surface of the scene. For the “w/o ray-splat intersection” case, we implement a basic 3D Gaussian splatting approach by adapting the projection calculation method specifically for panoramic maps. The integration of periodic vibration properties further improves GS-LiDAR’s capability in handling dynamic elements. Regularization terms, including median depth loss, depth distortion loss, and chamfer distance loss, contribute to the improved quality of the simulated LiDAR point clouds. Additionally, the ray-drop refinement technique improves the accuracy of the ray-drop mask, resulting in substantial gains in the metrics for simulated depth and intensity.

### v2

@@ Abstract
Figure 1: GS-LiDAR achieves superior LiDAR simulation quality for novel view synthesis while maintaining fast training and rendering speed.
@@ 3.4 LiDAR rendering
Figure 5: Comparison of 3D LiDAR point cloud. GS-LiDAR produces a more cohesive LiDAR point cloud compared to LiDAR-NeRF (Tao et al., 2023) and LiDAR4D (Zheng et al., 2024).
@@ 4.1 Experiment setup
We conduct extensive experiments on both dynamic and static scenes using the KITTI-360 (Liao et al., 2022) and nuScenes (Caesar et al., 2020) datasets, with the dynamic scenes featuring a significant number of moving vehicles. The KITTI-360 dataset employs a 64-beam LiDAR with a vertical field of view (FOV) of 26.4 degrees and an acquisition frequency of 10Hz. Following LiDAR4D (Zheng et al., 2024), we select 51 consecutive frames as a single scene and hold out 4 samples at 10-frame intervals for novel view synthesis (NVS) evaluation. For the nuScenes dataset, the LiDAR system uses 32 beams with a 40-degree vertical FOV and a 20Hz acquisition frequency. To ensure consistency with KITTI-360, we maintain a sampling frequency of 10Hz. We also provide the results on the Waymo (Sun et al., 2020) dataset in Appendix A.2.
Figure 6: Comparison of the rendered depth map with competitors.
@@ 4.2 Evaluation on static scenes
Figure 8: Comparison of w/ and w/o ray-splat intersection. A more detailed discussion is provided in Appendix A.1.
@@ 4.4 Ablation study
We provide quantitative ablation studies on various components of GS-LiDAR in Table 4. As shown in Figure 8, we find that the use of ray-splat intersection enhances the capability of GS-LiDAR in modeling the surface of the scene, and a more detailed discussion is provided in Appendix A.1. For the “w/o ray-splat intersection” case, we implement a basic 3D Gaussian splatting approach by adapting the projection calculation method specifically for panoramic maps. The integration of periodic vibration properties further improves GS-LiDAR’s capability in handling dynamic elements. Regularization terms, including median depth loss, depth distortion loss, and chamfer distance loss, contribute to the improved quality of the simulated LiDAR point clouds. Additionally, the ray-drop refinement technique improves the accuracy of the ray-drop mask, resulting in substantial gains in the metrics for simulated depth and intensity.
@@ Acknowledgment
Appendix A Appendix
A.1 Ray-splat intersection
Although 3DGS performs well in camera image rendering, it is not suitable for panoramic LiDAR rendering. This limitation arises from the complexity of transforming 3D spatial points into LiDAR range map, which makes it challenging for the Jacobian matrix of the transformation to accurately approximate the effects of such a transformation. Given a point $(x,y,z)$ in the camera space, its corresponding radian angles $(\phi,\theta)$ can be calculated as:
$\begin{pmatrix}\phi\\ \theta\end{pmatrix}=\begin{pmatrix}atan2(x,z)\\ atan2(\sqrt{x^{2}+z^{2}},-y)\end{pmatrix}$ (22)
By combining this with Equation 8 from the paper, we can derive the projection of this point onto the LiDAR range map, resulting in its pixel coordinates $(\xi,\eta)$ as:
$\begin{pmatrix}\xi\\ \eta\end{pmatrix}=\begin{pmatrix}\frac{W}{2\pi}atan2(x,z)+\frac{W}{2}\\ H\frac{atan2(\sqrt{x^{2}+z^{2}},-y)-\mathrm{VFOV}_{\min}}{(\mathrm{VFOV}_{\max}-\mathrm{VFOV}_{\min})}\end{pmatrix}$ (23)
Thus, the Jacobian matrix can be computed as follows:
$\mathrm{J}=\begin{pmatrix}\tilde{W}\frac{z}{x^{2}+z^{2}}&0&-\tilde{W}\frac{x}{x^{2}+z^{2}}\\ -\tilde{H}\frac{xy}{\sqrt{x^{2}+z^{2}}(x^{2}+y^{2}+z^{2})}&\tilde{H}\frac{\sqrt{x^{2}+z^{2}}}{(x^{2}+y^{2}+z^{2})}&-\tilde{H}\frac{yz}{\sqrt{x^{2}+z^{2}}(x^{2}+y^{2}+z^{2})}\end{pmatrix}$ (24)
where $\tilde{W}=\frac{W}{2\pi}$ and $\tilde{H}=\frac{H}{\mathrm{VFOV}_{\max}-\mathrm{VFOV}_{\min}}$.
We tile the plane with 2D and 3D Gaussian primitives of different colors and render their color and depth. As illustrated in Figure 9, we can observe that for 3D Gaussian primitives, due to the poor approximation of the Jacobian for the projection transformation, the originally tightly arranged points become scattered, which also leads to the discontinuity in depth shown in Figure 8 of the paper. In contrast, the 2D Gaussian primitives, being based on ray intersection without approximation, exhibit the desired effect.
Figure 9: Ray-splat intersection v.s. 3D Gaussian splatting.
A.2 Experiments on Waymo dataset
We additionally conducted experiments on sequences selected by PVG (Chen et al., 2023a) from the Waymo (Sun et al., 2020) dataset. Since methods like LiDAR4D (Zheng et al., 2024) have not been implemented on the Waymo dataset, we only report our own results in Table 5. The experimental results are consistent with those of KITTI-360 and nuScenes across the evaluation metrics, showing the scalability and generalization ability of our method. We have also additionally visualized LiDAR point clouds from the Waymo dataset in Figure 10 and Figure 11, showcasing the generalization ability of our approach again.
Table 5: Metrics on Waymo dataset.
Method Point Cloud Depth Intensity
[…]
Figure 10: Comparison of the rendered depth and intensity map with ground truth. The first two rows depict the depth maps, while the subsequent two rows illustrate the intensity maps.
Figure 11: The visualization of LiDAR point clouds in three-dimensional space.